Task 1


Estimate how much pressure you exert on the ground when standing up. Carefully account for the measurement data that you use when solving the problem
1
Expert's answer
2020-07-10T10:30:03-0400

Suppose during the measurements we obtain, that the area of one foot is "A_1 = 120 cm^2 = 0.012m^2". Then for both will be:


"A = 2A_1 = 2\cdot 0.012m^2 = 0.024m^2"

Also, suppose, the mass is "m =60 kg".

By definition, the preassure is given by:


"P = \dfrac{N}{A} = \dfrac{mg}{A}"

where "N" is the weight of body and "A" is the area of feet.

Thus:


"P = \dfrac{mg}{A} = \dfrac{60\cdot 9.81}{0.024} \approx 24.5\cdot 10^{3}Pa = 24.5\space kPa"

Answer. "P = \dfrac{mg}{A} =24.5 kPa".